OpenStudy (anonymous):

what is the equation of a line passing through (2,-1) and parallel to the line represented by the equation y=2x+1?

OpenStudy (andrefontex):

i think is this: y = k (2x+1) but -1 = k (2(2) +1) k = -1/5 then , y = -1/5 (2x+1)

OpenStudy (anonymous):

This should be in the mathematics group.